OpenStudy (anonymous):

Sasha solved an equation. Her work is shown below. Explain Sasha's error. Then, show the correct way to solve. h/7 = 6 h/7 + 7 = 6 + 7 h=13

OpenStudy (kidrah69):

\[(6)/7 = 6\] \[(13)/7 + 7 = 6 + 7\] well to double check her answer plug in 13 instead of h

OpenStudy (kidrah69):

then see where she made the mistake

OpenStudy (anonymous):

Thanks! @kidrah69

OpenStudy (anonymous):

7*(h/7) = 7*6 h = 42

OpenStudy (kidrah69):

After that you need to show the correct way to solve it we use the reciprocal so what @robtobey did

OpenStudy (anonymous):

Thanks @kidrah69 @robtobey

OpenStudy (anonymous):

Your welcome.